how do you graph the equation y=x^2-6x+6??? plllzz help....thannk u very much ^^

this is from a midyear exam in january 2004 in bronx science

Answer by stanbon(75887)

how do you graph the equation y=x^2-6x+6
---------
One way: You could plot points by giving values for
x and solving for y.
-------
Another way:
You might know the vertex is at x=-b/2a=6/2=3
Then find f(3)=9-18+6=-3
That gives you the vertex at (3,-3)
You can see that f(0)=6; that gives you the y-intercept.
You could find the x-intercepts by using the quadratic formula:
x=[6+-sqrt(36-4*6)]/2
x=[6+-sqrt(12)]/2
x= 3+sqrt3 or x=3-sqrt3
Cheers,
Stan H.